# Debate, Deliberate, Decide (D3): A Cost-Aware Adversarial Framework for Reliable and Interpretable LLM Evaluation

### v3

@@ 6.1 LLM-as-a-Judge
The concept of using a powerful LLM to evaluate the outputs of other models has become a cornerstone of modern LLM assessment Brown et al. (2020). This paradigm was popularized by studies demonstrating that models like GPT-4 can achieve high agreement with human preferences, offering a scalable alternative to manual annotation Zheng et al. (2023); Kim et al. (2024). Benchmarks such as MT-Bench and Chatbot Arena are built upon this principle, collecting pairwise comparisons to establish leaderboards of model performance Zheng et al. (2023). While foundational, this single-judge approach is the primary target for the improvements D3 aims to deliver, specifically addressing its documented vulnerabilities to cognitive and positional biases Wang et al. (2023); Mehrabi et al. (2022).
@@ 6.2 Multi-Agent Debate for Evaluation
To overcome the limitations of a single judge, researchers have explored multi-agent systems that simulate collaborative human evaluation processes Hong et al. (2024). ChatEval was a pioneering framework in this area, introducing the idea of a “referee team” of LLM agents with diverse personas that debate to reach a consensus Chan et al. (2024). This work established the potential of multi-agent debate to improve correlation with human judgments. PRD (Peer Rank and Discussion) further refined this by focusing on peer-review mechanisms to specifically target and reduce self-enhancement and positional biases Li et al. (2024); Wang et al. (2023). D3 builds directly on the insights from these systems. It incorporates the persona-driven approach of ChatEval and the bias-mitigation goals of PRD but introduces a more formalized, courtroom-inspired structure with distinct Advocate, Judge, and Juror roles. More importantly, D3’s primary novel contributions the explicit management of the cost-accuracy trade-off via its dual protocols and budgeted stopping rule, and its extensive, multi-benchmark empirical validation address key practical and scientific gaps left by these earlier frameworks.
@@ 6.3 Specialized Evaluator Models
An alternative to dynamic, prompt-based multi-agent systems is the development of specialized LLMs that are fine-tuned specifically for the task of evaluation Bai et al. (2022); Christiano et al. (2023); Ziegler et al. (2020). PandaLM is a prominent open-source example, a model trained on human preference data to act as a reliable and reproducible judge Wang et al. (2024). Similarly, Prometheus is another open-source LLM designed to replicate GPT-4’s evaluation capabilities when provided with appropriate criteria Kim et al. (2024). These models offer the advantage of being smaller, faster, and potentially more consistent for a specific evaluation task. However, they may lack the flexibility of prompt-based systems to adapt to novel tasks or evaluation criteria without retraining. In our experiments, PandaLM serves as a strong non-debate baseline, allowing us to demonstrate that the dynamic, deliberative process of D3 can outperform even a highly specialized, fine-tuned model.
@@ 6.4 Automated Evaluation Benchmarks
The empirical validation in this paper relies on the critical work done by the community in developing high-quality benchmarks Liang et al. (2023). MT-Bench provides a standardized testbed for general conversational abilities Zheng et al. (2023). AlignBench fills a crucial gap by offering a comprehensive, multi-dimensional benchmark for evaluating alignment in Chinese, which we adapt for our study Liu et al. (2024). Finally, AUTO-J introduces a large-scale dataset covering 58 real-world scenarios, with judgments provided by GPT-4, enabling evaluation across a wide spectrum of practical applications. Traditional metrics like BLEU Papineni et al. (2002), ROUGE Lin (2004), and their critiques Callison-Burch et al. (2006); Mathur et al. (2020) have informed the development of these modern benchmarks, while human evaluation standards Howcroft et al. (2020); Celikyilmaz et al. (2021) and reading comprehension benchmarks like SQuAD Rajpurkar et al. (2016) have shaped evaluation methodology. By validating D3 across these diverse and challenging benchmarks, this work ensures that its demonstrated performance is robust and generalizable, supporting the broader goal of building general language assistants aligned with human values Askell et al. (2021).
7 Discussion
Our results show that structured, multi-agent frameworks such as D3 provide a more reliable and interpretable alternative to single-judge evaluations. By explicitly modeling the cost–accuracy frontier, D3 reframes evaluation as a trade-off rather than a single metric, helping practitioners select appropriate protocols for different development stages.
This work thus highlights the need to treat cost as a core dimension in evaluation research. Beyond improving accuracy, future work should aim to make high-fidelity evaluations more accessible through distillation, automated role generation, and more efficient coordination mechanisms among agents.
@@ 8 Conclusion
In this paper, we introduced Debate, Deliberate, Decide (D3), a multi-agent framework designed to address critical gaps in LLM evaluation. Faced with the biases of single-LLM judges and the scalability limits of human annotation, D3 provides a structured, adversarial process that leverages role-specialized agents to produce more reliable assessments. Through a comprehensive empirical study, we demonstrated that D3 significantly outperforms strong baselines in agreement with human judgments. Crucially, our work moves beyond accuracy alone, presenting a systematic methodology for auditing biases and introducing the cost-accuracy frontier as a vital tool. By delivering a framework that is empirically robust, systematically audited, and practically cost-aware, D3 represents a significant step towards the scalable, reliable, and trustworthy evaluation of AI systems.
@@ 9 Limitations
Computational Cost. Although D3 is designed to be cost-aware, it remains substantially more expensive than single-judge evaluation. The D3-MORE protocol requires roughly four times the tokens of a single-judge setup, and D3-SAMRE can consume even more. This additional cost may be justified for high-stakes assessments or final validation but can be prohibitive for early-stage, iterative testing. The cost–accuracy frontier in Section 5.3 aims to make this trade-off explicit.
Persona Design and Fairness. The use of diverse juror personas, while empirically beneficial, risks reinforcing social or cultural stereotypes. Responsible persona construction demands careful auditing to ensure diversity does not compromise fairness or neutrality. Developing principled, automated methods for persona generation remains an open challenge in human–AI interaction and responsible AI research.

### v4

@@ 6.1 LLM-as-a-Judge
Using a strong language model to evaluate the outputs of other models has become a central paradigm in modern LLM assessment Brown et al. (2020). Prior work has shown that models such as GPT-4 can achieve high agreement with human preferences, enabling scalable evaluation without extensive manual annotation Zheng et al. (2023); Kim et al. (2024). Widely used benchmarks including MT-Bench and Chatbot Arena operationalize this paradigm through pairwise comparisons judged by a single LLM Zheng et al. (2023).
Despite its practicality, the single-judge paradigm is known to suffer from systematic failure modes, including positional bias, verbosity bias, and preference leakage from prompt structure Wang et al. (2023); Mehrabi et al. (2022). Recent work has further shown that LLM judges may inadvertently encode information about the evaluated model or task context, leading to biased or unstable evaluations Li et al. (2025). D3 targets this setting directly, treating single-judge evaluation as a strong but brittle baseline and explicitly addressing its bias and robustness limitations through structured deliberation and anonymization.
@@ 6.2 Multi-Agent Debate for Evaluation
Motivated by the limitations of single-judge evaluation, several works have explored multi-agent debate as a mechanism for approximating collective human judgment Hong et al. (2024). ChatEval Chan et al. (2024) introduced the idea of a referee team composed of LLM agents with distinct personas who debate before reaching a final verdict, demonstrating improved correlation with human judgments. PRD (Peer Rank and Discussion) Li et al. (2024) further refined this approach by emphasizing peer-review-style discussion to mitigate self-enhancement and positional biases Wang et al. (2023).
More recent work has expanded this line of research. KIEval (Yu et al., 2024) addresses data contamination through multi-round interactive dialogues grounded in domain knowledge, enabling evaluation of genuine model understanding over memorization. M-MAD (Feng et al., 2025) decomposes machine translation evaluation into four orthogonal error dimensions and uses multi-agent debate within each dimension to refine error detection and severity assessment, improving segment-level accuracy. In parallel, Zhang et al. (2025) demonstrate through comprehensive benchmarking that homogeneous multi-agent debate rarely outperforms simple baselines despite higher inference costs, arguing that model heterogeneity—not unconstrained deliberation—is essential for effective collaborative reasoning.
D3 builds on these insights while addressing their practical limitations. Like ChatEval and PRD, it leverages role specialization and debate to reduce bias; however, it introduces a courtroom-inspired structure with explicitly defined Advocate, Judge, and Juror roles. Crucially, D3 differs from prior debate frameworks by explicitly managing the cost–accuracy trade-off through dual evaluation protocols and a budgeted stopping rule. This design directly responds to recent critiques of debate-based evaluation, enabling controlled deliberation rather than assuming that more agents or longer discussions are always beneficial.
@@ 6.3 Specialized Evaluator Models
An alternative to prompt-based or debate-driven evaluation is to train specialized models that act as judges Christiano et al. (2023); Ziegler et al. (2020); Bai et al. (2022). PandaLM exemplifies this approach as an open-source evaluator trained on human preference data to produce consistent and reproducible judgments Wang et al. (2024). Similarly, Prometheus aims to replicate GPT-4-level evaluation when conditioned on explicit evaluation criteria Kim et al. (2024).
Specialized evaluators offer efficiency and stability advantages but are inherently limited by their training distribution and require retraining to adapt to new tasks or criteria. Recent analyses in 2025 have further highlighted their vulnerability to preference leakage and overfitting to benchmark-specific annotation styles Li et al. (2025); Marioriyad et al. (2025). In our experiments, PandaLM serves as a strong non-debate baseline, allowing us to isolate the contribution of D3’s deliberative process and show that structured, budget-aware debate can outperform even a highly optimized evaluator model.
@@ 6.4 Automated Evaluation Benchmarks
This work relies on the substantial community effort devoted to building robust evaluation benchmarks Liang et al. (2023). MT-Bench provides a standardized benchmark for general conversational ability Zheng et al. (2023), while AlignBench offers a comprehensive multidimensional evaluation of alignment in Chinese, which we adapt for our study Liu et al. (2024). AUTO-J Li et al. (2023) introduces large-scale evaluation across 58 real-world scenarios using GPT-4 judgments, enabling broad empirical coverage.
These benchmarks build on earlier work in automated metrics such as BLEU and ROUGE Papineni et al. (2002); Lin (2004), as well as long-standing critiques of their correlation with human judgment Callison-Burch et al. (2006); Mathur et al. (2020). Human evaluation protocols Howcroft et al. (2020); Celikyilmaz et al. (2021) and task-specific benchmarks such as SQuAD Rajpurkar et al. (2016) have further shaped modern evaluation methodology. By validating D3 across multiple benchmarks with differing assumptions and failure modes, we aim to demonstrate that its improvements are robust rather than benchmark-specific, contributing to the broader goal of reliable and scalable alignment evaluation Askell et al. (2021).
@@ 7 Conclusion
D3 addresses critical gaps in LLM evaluation through structured, cost-aware multi-agent debate. Across MT-Bench, AlignBench, and AUTO-J, D3 outperforms baselines in accuracy, positional consistency, and self-enhancement robustness. D3-MORE provides efficiency comparable to ChatEval while achieving +8.3% higher accuracy than PRD. D3-SAMRE with budgeted stopping achieves highest accuracy at 4.65× single-judge cost, with 58% of debates converging by round 2. This cost-accuracy frontier enables practitioners to select protocols matching their constraints. By combining role specialization, systematic bias auditing, and explicit cost-awareness, D3 advances reliable and scalable LLM evaluation. Future work could explore automated role generation and distillation methods to further democratize access to high-fidelity evaluation.
@@ 8 Limitations
Computational Cost. Although D3 is designed to be cost-aware, it remains more expensive than single-judge evaluation. The D3-MORE protocol requires roughly four times the tokens of a single-judge setup, and D3-SAMRE can consume even more. This additional cost may be justified for high-stakes assessments or final validation but can be prohibitive for early-stage, iterative testing. The cost–accuracy frontier in Section 5.4 aims to make this trade-off explicit.
Persona Design and Fairness. Our diverse juror personas provide empirical benefits (Section 5.1) but require responsible fairness management. Personas are role-based (ethics professor, business owner, social worker, environmental activist, tech entrepreneur) rather than demographic, reducing stereotype risks. All jurors use identical backbone LLM; personas only influence instructional framing. Section 5.1 validates robustness through controlled ablations: persona effects are stable, complementary ($89.5\%$ cross-persona agreement), and free of demographic bias (effect sizes <0.04<0.04). Open challenges remain: cross-cultural generalization in non-Western contexts, automated persona generation for scalability, and fairness audits across protected characteristics of response subjects. Future work should conduct targeted fairness studies to ensure D3 does not systematically advantage particular demographic groups.
@@
9 Acknowledgements
We thank Amir Abdullah, Philip Quirke and Michael Lan for their invaluable feedback on the project. We also appreciate Antia Garcia Casal for her help with figures and for shaping the overall style of our visualizations.
@@ 6. Effectiveness in addressing opponent’s points
Appendix G Persona Sensitivity Analysis: Full Ablation
G.1 Systematic Persona Pool and Robustness Analysis
We constructed a diverse pool of 50 personas spanning law, medicine, education, technology, ethics, business, social work, risk analysis, and compliance. We then conducted 10 independent experiments on MT-Bench (120 randomly selected questions), each using a different random 5-persona subset. All personas were instantiated with GPT-4-Turbo as the backbone; personas influenced only instructional prompts, not underlying model capabilities. Table 4 summarizes results:
Table 4: Persona Robustness Analysis: Ablation Across Random 5-Persona Subsets (MT-Bench, 120 questions)
Configuration Accuracy (%)
Random 5-persona subsets (avg of 10) $85.9\pm 0.7$
Curated 5-persona set $86.1$
Generic jurors (no personas) $82.7\pm 1.1$
(1) Persona Conditioning Effect:
Persona-guided evaluation outperforms generic jurors by $3.2\%$ (85.9% vs. 82.7%, $p<0.01$). The persona effect holds robustly across random subsets, indicating diversity, not specific persona identity, drives performance.
(2) Robustness Across Subsets:
Tight variance ($0.7\%$ std. dev.) indicates D3’s performance is not brittle to persona composition; the framework generalizes well to unseen persona combinations.
(3) Curated vs. Random:
Curated personas ($86.1\%$) achieve marginally higher accuracy than the random average ($85.9\%$, $0.2\%$). This validates that deliberate design adds value without harming generalization.
G.2 Design Rationale: Persona Selection and Value Lenses
Our five curated personas were selected to provide complementary professional perspectives and value lenses. Each persona is anchored in domain expertise rather than demographic attributes, reducing stereotype risks:
Framework Evaluator Accuracy (%) Relative Gain
Single Judge Llama-3-70B 68.3 baseline
ChatEval Llama-3-70B 73.1 +4.8%
PRD Llama-3-70B 71.5 +3.2%
D3-MORE Llama-3-70B 73.9 +5.6%
D3-MORE GPT-4-Turbo 85.1 +12.6%
Table 5: MT-Bench accuracy using open-source and proprietary evaluators.
  • Ethics Professor: Ethical principles, long-term societal impact
  • Environmental Activist: Collective welfare, ecological responsibility
  • Business Owner: Practical feasibility, roi, trade-offs
  • Social Worker: Human-centered perspective, equity, vulnerable populations
  • Tech Entrepreneur: Innovation, scalability, technological progress
This composition ensures verdicts incorporate ethical grounding, distributional impact, economic viability, and technological feasibility. These roles are intentionally role-based rather than demographic; all jurors execute the same backbone model with identical base capabilities.
G.3 Cross-Persona Agreement and Complementarity
We computed pairwise vote agreement across the five curated personas. Average agreement is $89.5\%\pm 1.5\%$, with pairwise ranges $87{-}92\%$. High cross-persona agreement confirms personas are complementary rather than contradictory, reducing the risk of arbitrarily conflicting judgments while maintaining diverse perspectives.
G.4 SAMRE Debate Progression Across Question Types
The convergence behavior of iterative debate varies systematically across task complexity and reasoning demands. We analyze score gap trajectories across MT-Bench question categories to validate that the budgeted stopping rule terminates debates at their point of maximal signal without sacrificing verdict reliability.
Coding tasks exhibit the largest discriminative gaps, with peaks reaching 20 points by round 2 and maintaining plateau stability through round 5. This reflects clear correctness boundaries in code evaluation. Reasoning and ethics questions show moderate, steady gaps (8–11 points) with gentle convergence, indicating these subjective domains benefit from iterative refinement but stabilize predictably. Writing, roleplay, and math tasks display tighter gaps (0–6 points) and earlier plateau behavior, suggesting these domains reach verdict saturation with fewer debate rounds.
Across all categories, gap stabilization occurs by rounds 4–5, with the majority of verdicts determined by round 2. The final round markers (diamonds) per question type show that D3-SAMRE can terminate heterogeneously: early for low-variance tasks like math and writing, while extending slightly longer for nuanced reasoning and ethics evaluations. This task-sensitive convergence supports the empirical guidance in Section 5.5: use SAMRE for high-stakes and ethically complex evaluations, while MORE suffices for objective or well-separated task types.
Figure 5: Score gap trajectories across six MT-Bench question categories reveal differential convergence patterns in iterative adversarial debate. Coding tasks exhibit the largest gaps (peak 20 points) with stable plateaus, indicating clear discriminability. Reasoning and ethics tasks show moderate, steady gaps (8–11 points). Writing, roleplay, and math questions display tighter gaps (0–6 points) with earlier plateau behavior. Diamond markers indicate final verdict round per category.
Appendix H Open-Source Evaluator Results
To test whether D3’s performance gains depend on evaluator model quality, we evaluate all frameworks using Llama-3-70B Llama (2024) as the evaluator on 100 MT-Bench Zheng et al. (2023) questions. Table 5 reports accuracy and relative gains over a single-judge baseline.
These results show that D3’s improvements persist when using a fully open-source evaluator, indicating that the gains stem from the evaluation architecture rather than reliance on proprietary judge models.